OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±2) and foci at (0, ±11).

OpenStudy (anonymous):

The hyperbola is vertical so the equation is gonna be in the form of: \[(y^2/a^2) -(x^2/b^2) = 1\]

OpenStudy (anonymous):

we know a (2) and we know c (11)

OpenStudy (anonymous):

so you can use the pythagorean theorem to find b: a^2 + b^2 = c^2 2^2 + b^2 = 11^2 4 + b^2 = 121 b^2 = 117

OpenStudy (anonymous):

so the equation would be: (y^2/4) - (x^2/117) = 1